When equal amounts are invested in each of three accounts paying 7% 8% and 10.5% interest, one year's combined interest income is $1,109.25. How much is invested in each account?
Answer by richwmiller(17219)

0.07*z+0.08*z+0.105*z=1109.25
x=y
y=z
.255*z=1109.25
z=1109.25/0.255
x=4350
y=4350
z=4350
x+y+z=13050
0.07*4350+0.08*4350+0.105*4350=1109.25